OpenStudy (anonymous):

The perimeter of a basketball court is 90 meters and the length is 6 meters longer than twice the width . What are the length and width.

Directrix (directrix):

Let w = width of the court Then, l = 2w + 6 2w + 2(2w + 6) = 90 2w + 4w + 12 = 90 6w + 12 = 90 6w = 78 w = 13 l = 2*13 + 6 = 32 Length = 32 and Width = 13

OpenStudy (campbell_st):

 Now form an equation to represent the perimeter... and then solve for x